# HUG: Hierarchical Urban Gaussian Splatting with Block-Based Reconstruction

### v1

@@ 3.4 Local Scene Fusion and Rendering
We are the first to introduce neural octree Gaussians to urban scene reconstruction tasks. After training each block in parallel, we refilter the anchors within each block. We apply a a non-linear transformation similar to that used in CityGS[17] to map all anchor centers to the range [-2, 2]. The bounding box obtained from the block index is then used to refilter the anchors. Since each anchor has 10 neural Gaussians, and the offsets can sometimes be quite large, some anchors may be located inside the bounding box while their neural Gaussians are outside, or vice versa. Therefore, the positions of the neural Gaussians are also considered during the refiltering process. An anchor with more than 5 Gaussians inside the bounding box will be retained in the block, while anchors with fewer than 2 Gaussians inside the bounding box will be discarded, even if the anchor itself is within the block. Refiltered anchors of all blocks are then merged to form a global scene with an additional block index property.
@@
During rendering, we first load the global scene into the GPU. Although each block has a separate MLP network, we load all MLP networks to the GPU and use the block index to select the corresponding MLP network for each anchor. These MLPs are small and require very little GPU memory. The MLP network is used to calculate the neural properties of the Gaussians, which are then applied to render the scene using differentiable rasterization from 3DGS [12]. This approach enables global scene rendering and viewing without necessitating LOD file loading.
Figure 4: Qualitative comparison with state-of-the-art methods on synthetic and real-world datasets. We select CityGaussian[17] as our sole baseline because it is the only reproducible method evaluated on both synthetic and real-world datasets. Our method reconstructs finer details and produces fewer artifacts compared to baseline.
@@ 4.1 Implementation Details
We built our method upon Octree-GS[26] for implementation. We modified the rendering codes and SIBR viewer[4] to support LOD rendering of neural Gaussians generated by MLPs from multiple blocks, enabling real-time rendering of large-scale urban scenes. All source code, datasets, and preprocessing scripts will be made publicly available upon publication.
Our training process consists of two steps. First, we partition the data using the sparse reconstruction results from COLMAP[27], which enables parallel scene reconstruction across multiple GPUs. Specifically, we divide the bounding box into several blocks along the coordinate axes, tailored to the shape of the scene. The sparse points within each block are then projected into image space, and an SVM[6] is used to fit the boundary between inner and outer points after a tenfold downsampling. Additionally, we assign ten times the weight to points within the block, obtaining a mask for the inner and outer parts of the scene in image space. Finally, for each block, we select cameras with a visible area greater than $\tau_{p}=10\%$ of the image area to participate in the training of the block’s scene.
In the subsequent phase, we performed block-wise reconstruction initializing from the sparse points seen by selected cameras. These sparse points are preprocessed similar to OctreeGS[26]. We first train for 40,000 iterations as initiation. In this stage, we split or prune anchors every 50 iteration until iteration 30,000, the gradient threshold degeneration factor ${\eta}$ is 0.8, and the interval $M$ is 5,000. The Manual splitting ratio $\tau_{s}$ is 0.001. When $mean(L_{e})>1$, we add 1 to the octree level $K$, and all $L_{e}$ aree divided by 50. For pruning anchors, the visibility count ${\epsilon_{c}}$ is 5, and accumulated gradient threshold ${\epsilon_{g}}$ is set as 0.000002. We render anchros of all levels after iteration 10,000. Then our visibility mask and weighted hierarchical supervision are applied for other 40,000 iterations refinement separately. The weight $\lambda$ between L1 and SSIM is 0.2, the weight $\theta$ is set as 0.02. And the opacity mask is $m^{o}_{L}>0.5$. All experiments were conducted on a machine with 8 Nvidia A6000 48GB GPUs and 2 Intel Xeon 6330 CPUs.
@@ 4.2 Evaluation
We evaluated our method on four diverse scenes with varying scales and environments. We utilized a synthetic SMall City scene from the city-scale dataset, MatrixCity[14], which is 2.7 $km^{2}$ in size. Unlike previous approaches focusing on partial areas[29], we reconstructed the entire city to assess our method’s performance comprehensively. In addition to synthetic data, we conducted experiments on real-world datasets, including Residence, Rubble, and Building from Mega-NeRF[33]. Consistent with prior work[33, 37, 36, 17, 5, 26, 18], we reduced the image resolutions to 1.6k to balance detail and processing time.
For quantitative evaluation, we employed standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS)[35].
We compared our method against several state-of-the-art baselines, including Mega-NeRF[33], 3DGS[12], CityGaussian[17], and OctreeGS[26].
@@ 4.3 Results Analysis
Table 1: Quantitative comparison of our method with previous approaches. The best scores in each column are highlighted with red background, second best with orange, and third best with yellow. The symbol ’-’ denotes experiments that are difficult to reproduce.
MatrixCity Residence Rubble Building
@@
MegaNeRF[33] - - - 0.628 22.08 0.489 0.553 24.06 0.516 0.547 20.93 0.504
3DGS[12] 0.735 23.67 0.384 0.791 21.44 0.236 0.777 25.47 0.277 0.720 20.46 0.305
Octree-GS[26] 0.814 26.41 0.282 - - - - - - - - -
CityGS[17] 0.865 27.46 0.204 0.813 22.00 0.211 0.813 25.77 0.228 0.778 21.55 0.246
Ours 0.883 28.02 0.142 0.804 22.18 0.231 0.796 25.80 0.247 0.761 21.91 0.276
Quantitative Comparisons
Table 1 shows the mathematical results, our method achieves SOTA performance compared with novel NeRF-based and Gaussian-based methods on the MatrixCity[14] dataset. Our method outperform OctreeGS[26] over 1.5 PSNR, which shows the effectiveness of block-based reconstruction and global fusion in our method. We also overwhelm the SOTA method CityGS[29] on this dataset, this proves the effective of our combination of hierarchical representation and neural Gaussian. Our method achieves competitive results on 3 scenes from real world dataset. We again performs better than CityGS, which has similar partition strategy with our work, showing the in-block efficiency of our hierarchical representation and novel fusion technology. We averaged fewer cameras per block than CityGS as shown in the supplement, yet our approach still outperformed theirs.
@@
Figure 4 showcases qualitative comparisons between our method and the SOTA baseline CityGS[17] on three scenes. Our method produces sharper and more detailed reconstructions, accurately capturing fine textures and structural details with fewer artifacts.
In particular, our reconstructions exhibit better preservation of intricate features such as building facades, textures, and edges. The car we rendered are even clearer than the ground truth. We better restored the lighting intensity of the lawn, showing the effectiveness of our weighted hierarchical supervision. The visual fidelity is consistent across different scenes, highlighting the robustness of our approach.
Efficiency
We are the first to introduce a hierarchical neural Gaussian pipeline to urban scene reconstruction task, our main purpose is to allow partitioned parallel training and render the whole urban scene with low memory consumption, we compare with CityGS[17] both in LOD mode and non-LOD mode. The CityGS-LOD mode uses Gaussian compression method LightGaussian[8] to compress Gaussians and reduce computational cost. We render Rubble scene with similar visual quality while gradually raising the view center. The results are shown in Figure 5. As the view height increases, both CityGS and CityGS-LOD have a increase in number of Guassians. But, with our hierarchical representation, the number of Gaussians initially increases and then decreases. Our rendering speed outperforms CityGS, though it is slightly slower than CityGS-LOD. We believe this discrepancy is due to differences in compression rates, and our method involves selecting anchors and predicting neural properties.
Figure 5: Rendered Gaussian count and FPS performance tested by increasing the camera height in the $Rubble$ scene.
4.4 Ablation
To thoroughly evaluate the contributions of each component in our proposed method, we conducted ablation study on the Building scene. Specifically, we examined the effects of the usage of the splitting and pruning tricks (S&P) mentions in Section 3.3, the hierarchical supervision (Hier.) proposed in Equation 8, and the block-wise focused training (Focused) strategy mentioned in Section 3.2.
Table 2: Ablation on splitting and pruning tricks, hierarchical supervision and focused training.
@@ 4.4 Ablation
w/o S&P 0.695 20.78 0.355
w/o Hier. 0.731 21.12 0.309
w/o Focused 0.749 21.44 0.285
Full model 0.761 21.91 0.276
Splitting and Pruning Strategy
Splitting is an important part in Gaussian-based methods, insufficiency of splitting will lead to the degeneration of the rendered image. In our work, we proposed some novel splitting and pruning tricks specifically designed for the hierarchical neural Gaussian urban scene reconstruction task. These tricks can effectively split anchors in detail-rich areas and prune anchors in tricky positions. Our methods is able to increase the LOD levels of the scene automatically. All these functions promise our model higher representation ability and better performance as shown in Table 2.
Hierarchical Supervision
In our work, we proposed a novel weighted hierarchical supervision constrain for training layered representations in Equation 8. This supervision compensates for the deficiencies of the anchor selection in OctreeGS[26]. By rendering anchors belonging to different anchors separately, and metric them with different weights, we can effectively learn the hierarchical representations. This supervision effectively guides the model to learn the local and global details of the urban scene, which is crucial for the high quality reconstruction.
Focused Training Strategy
Our method employs a block-wise focused training strategy to refine specific regions of the scene by calculating the loss only at locations where the visibility mask equals 1. As shown in Table 2, the block-wise focused training strategy significantly improves reconstruction quality. We believe this is because our strategy reduces the interference from areas with less coverage outside the block on the training within the block.
@@ 5 Conclusion
In this paper, we present HUG, a novel Hierarchical Urban Gaussian Splatting method designed to efficiently reconstruct and render large-scale urban scenes. The experimental results, demonstrating State-of-the-Art performance on public benchmarks. The efficiency of our HUG for scene reconstruction and rendering sets new standards for urban scene processing.
Despite our approach’s advantages, a few limitations should be addressed in future work. First, our method does not account for dynamic objects or lighting inconsistencies that may exist in real-world large-scale scene datasets. This can be improved by incorporating lighting decoupling and appearance encoding. Second, the data partitioning process relies on the sparse point cloud generated by COLMAP[27], which may lead to issues in low-texture areas that lack sufficient sparse points. Combining pose-free Gaussian methods[11] may help it.
@@
We provide a more detailed discussion on the necessity of visibility mask, data partitioning results, the demonstration of issues with unoptimized anchor points, and more qualitative comparisons.
6 The Necessity of Visibility Mask
Figure 6: Illustration of the importance of our visibility mask. (a). By employing the visibility mask, our approach efficiently excludes irrelevant regions outside the block, concentrating computational resources on the in-block area. (b). The upper subfigure illustrates the typical approach adopted by most existing methods, where reconstructing a block often entails processing extensive extraneous areas, which are subsequently cut during the fusion step. In contrast, the lower subfigure demonstrates our method, which prioritizes the in-block regions, leading to a more focused reconstruction process.
One of our key observations is that areas outside the reconstruction block do not need to be rebuilt. Existing methods often rebuild extensive regions outside the block, which will be cut during the fusion step. This is depicted as gray areas in Figure 6. By applying the visibility mask, we effectively exclude irrelevant regions from reconstruction, concentrating computational resources on enhancing reconstruction quality within the block. This strategy enables high-quality reconstruction both within the interior of the blocks and at their boundaries without the need to expand the blocks themselves. As a result, we optimize computational efficiency and hardware resource utilization while maintaining control over block sizes and avoiding memory overflow. To achieve this, we apply a non-linear classification method to segment the visible regions from the projected sparse point cloud.
@@ 7 Data Partitioning Results
Figure 7: Comparison of the data partition results between our method and CityGS[17] across four datasets. The radial coordinate represents the number of training images of each block, while each vertex corresponds to a specific block. It is evident that our method assigns fewer images per block.
To associate all training images with respective blocks, we calculate the total visible area in our generated visibility masks. A threshold of $\tau_{p}=10\%$ is then applied to assign each training image to specific blocks. In CityGS[17], a coarse global 3DGS[12] model must first be constructed, after which all images are rendered multiple times corresponding to the number of blocks. As a result, this partitioning process is time-intensive, typically taking hours to complete, e.g., 1.1 hours for the Rubble scene. In our method, the partitioning only takes 10 minitus, which is much faster. Moreover, our data partitioning strategy assigns fewer images to each block, reducing memory requirements during training and allowing more resources to be dedicated to reconstructing the in-block area. The number of training images per block is visualized in Figure 7, while the average number is presented in Table 3. We can see that we allocate fewer images to each block, while still achieving better rendering performance as demonstrated in Table 1.
@@
Figure 9: More qualitative comparisons on MatrixCity dataset.
Table 3: The average number of training images per block in four scenes.
Methods MatrixCity Rubble Building Residence
CityGS 438.7 462.0 280.0 697.4
Ours 337.3 379.6 199.8 335.2
8 Demonstration of Issues with Unoptimized Anchor Points
Figure 10: The unoptimized anchor problem. (a). The original rendered view. (b). As the perspective zooms in, some multicolored Gaussians appear. (c). As the perspective zooms in further, more multicolored Gaussians appear, while our method solves this problem. (d). Local zoom-in.
In Octree-GS[26], only anchors with a visibility count greater than a certain threshold are considered for rendering, optimization, and pruning. Additionally, the levels of the anchors are taken into account during training. These strategies enable hierarchical training, but they also introduce some issues. Some anchor points are insufficiently optimized, resulting in reconstruction errors. As shown in the upper row of Figure 10, the anchor points are insufficiently optimized, resulting in the rendering of distorted, multicolored Gaussians. This phenomenon becomes more pronounced as the perspective narrows. In our work, we address this issue through several strategies. First, we prune anchors with a visibility count below a certain threshold ${\epsilon_{c}=5}$. Second, we prune anchors with an accumulated gradient below a specified threshold ${\epsilon_{g}=0.000002}$. Third, we render all visible anchors regardless of their LOD level at certain stages of training. These approaches collectively help optimize the anchor points and reduce reconstruction errors as shown in the lower row in Figure 10.

### v2

@@ 3.4 Scene Fusion and Rendering
After completing training in all blocks, previous works, such as CityGS [17], typically refilter Gaussians whose center positions lie outside the block boundaries. However, we found that this approach can introduce artifacts near the boundaries. Specifically, even if an anchor’s center is outside the block, the neural Gaussians it generates often have centers that fall within the block.
To address this, we propose an innovative greedy vote-based anchor refiltering strategy. For any given anchor, it is retained if its center lies within the block. If the anchor’s center is outside the block but more than half of its generated neural Gaussians have centers within the block, the anchor is also preserved. Notably, our method follows a greedy approach, as it selectively retains anchors based on voting results but does not actively remove them.
@@
After refiltering, all blocks are stitched together. Methods like CityGS [17] can directly leverage the 3DGS [11] rasterization pipeline for global rendering. However, methods such as Octree-GS [26], which incorporate neural Gaussians, rely on small MLPs. These MLPs are trained separately for different neural networks and thus have distinct weights, making Octree-GS unsuitable for direct application to multi-block scenes. To overcome this limitation, we reimplement the rasterization process to support the loading of multiple MLPs. During rendering, all visible anchors are first selected, and their corresponding neural Gaussians are then generated in parallel using the MLP trained on the respective block. This multi-MLP rendering strategy strikes a balance between leveraging the efficiency of MLPs and accommodating their limited learning capacity, enabling seamless rendering across blocks.
Figure 4: Qualitative comparison with state-of-the-art methods on one synthetic and four real-world datasets. Red insets highlight patches that reveal notable visual differences between these methods.
@@ 4.1 Implementation Details
We built our method on the foundation of Octree-GS [26] for implementation. Our framework integrates innovative techniques, including visibility-based partitioning, hierarchical weighted image supervision, and a novel anchor splitting and pruning strategy. We modified the rendering code and the SIBR viewer [4] to support LOD rendering of neural Gaussians generated by multiple MLPs from different blocks, enabling real-time rendering of large-scale aerial urban scenes.
We first apply our proposed visibility-based partitioning method to divide the scene and views into multiple blocks. The point cloud threshold $\tau_{p}$ is setting as 800. We then perform block-wise scene optimization. The weights balancing components in the $l_{hier}$ are set as $\lambda=0.2$, $\gamma=0.8$, and $\theta=0.02$. For every $M=5000$ iterations training, anchors with visibility count lower than ${\epsilon_{c}}=5$ will be pruned, and gradient threshold $\epsilon_{g}=2e^{-6}$ will be decreased by $\eta=0.8$. The $\beta$ in dynamic anchor level transition is set as 4. And the opacity mask $m^{o}_{L}$ is obtained with pixels with opacity larger than 0.5. All experiments were conducted on a machine with 8 Nvidia A6000 48GB GPUs and 2 Intel Xeon 6330 CPUs.
@@ 4.2 Evaluation
We evaluated our method on five diverse scenes spanning different scales and environments. Our experiments included a synthetic Small City scene from the city-scale dataset MatrixCity [14], covering an area of 2.7 $km^{2}$. Beyond synthetic data, we tested our approach on real-world datasets, including Residence, Rubble, Building, and Sci-Art from Mega-NeRF [33]. Following prior works [33, 37, 36, 17, 5, 26, 18], we downsampled the four real-world datasets by a factor of 4 and reduced the image resolution to 1.6k for MatrixCity.
For quantitative evaluation, we employed standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS[35].
We compared our method against several state-of-the-art baselines, including Mega-NeRF [33], 3DGS [11], CityGS [17], OctreeGS [26], VastGS [15], and Hier-GS [12]. More details on the settings for each method are provided in the supplement.
@@ 4.3 Results Analysis
Quantitative Comparisons
Table 1 presents the quantitative results. Except for the PSNR metric in the Sci-Art scene, possibly due to the large number of distant views, where our results are lower than Mega-NeRF [33] but still outperform all GS-based methods. Across all other metrics, it surpasses all baseline methods. Notably, our method exceeds Octree-GS [26] by over 1.5 PSNR, demonstrating the effectiveness of our block-based reconstruction and other proposed innovations. Additionally, we achieve superior performance compared to CityGS [29] and VastGS [15], further validating the effectiveness of our hierarchical representation combined with neural Gaussians. Our method also outperforms Hier-GS [12], which we attribute to our specific design tailored for large-scale aerial scenarios. These results further underscore the efficacy of our visibility-based partitioning strategy.
MatrixCity Residence Rubble Building Sci-Art
@@
MegaNeRF [33] - - - 0.628 22.08 0.489 0.553 24.06 0.516 0.547 20.93 0.504 0.770 25.60 0.390
3DGS [11] 0.735 23.67 0.384 0.791 21.44 0.236 0.777 25.47 0.277 0.720 20.46 0.305 0.830 21.05 0.242
Octree-GS [26] 0.814 26.41 0.282 - - - - - - - - - - - -
Hier-GS [12] 0.842 26.67 0.251 0.776 20.24 0.221 0.765 22.28 0.257 0.733 20.04 0.262 0.828 19.74 0.207
VastGS [15] - - - 0.777 20.38 0.247 0.766 25.10 0.294 0.740 21.69 0.293 0.809 21.66 0.269
CityGS [17] 0.865 27.46 0.204 0.813 22.00 0.211 0.813 25.77 0.228 0.778 21.55 0.246 0.837 21.39 0.230
Ours 0.883 28.02 0.142 0.813 22.33 0.207 0.839 26.42 0.197 0.792 22.35 0.228 0.846 21.83 0.204
Table 1: Quantitative comparison of our method with state-of-the-art methods. The best scores in each column are highlighted with red background, second best with orange, and third best with yellow. The symbol ‘-’ denotes experiments that are difficult to reproduce.
@@
Figure 4 provides qualitative comparisons between our method and others across five diverse scenes. In the MatrixCity scene, our method excels at reconstructing intricate details, such as rooftop textures, facade steel structures, and street lamps, while preserving the fine structures with high fidelity. In the Residence scene, our approach captures the truck’s details more thoroughly than the baseline methods. In the Rubble scene, while all methods successfully reconstruct the red object, our method stands out by exhibiting fewer artifacts in the surrounding rubble piles. In the Building scene, both our method and Hier-GS [12] reconstruct the staircase steps accurately, but Hier-GS [12] suffers from noticeable color deviations. Finally, in the Sci-Art scene, our method successfully recovers a building nestled between two background structures, which other methods fail to reconstruct entirely.
Figure 5 also illustrates how our method effectively eliminates the unoptimized anchor issue in Octree-GS [26], which often becomes apparent when zooming in on the scene.
Figure 5: The unoptimized anchor issue of Octree-GS [26]. (a). A training view. (b). As the perspective zooms in, multicolored artifacts begin to appear in Octree-GS [26]. (c). Further zooming amplifies these artifacts, whereas our method effectively eliminates them. (d). Close-up view of the issue.
Partitioning and Optimization Efficiency
Partitioning large-scale scenes into smaller blocks followed by optimization has become a standard practice. Since these two stages are executed sequentially, the time consumption of each directly impacts the overall performance. Table 2 presents the time consumption of our method compared to related approaches. Our visibility-based partitioning is both simple and efficient, consistently completing in about one minute, whereas CityGS [17] requires over two hours for the MatrixCity dataset. Efficient optimization within each block is equally crucial. While most methods, including ours, maintain an optimization time of approximately one hour per block, Hier-GS [12] takes nearly two hours due to additional post-processing refinements applied to each block.
Scene MatrixCity Rubble Residence Building Sci-Art
Task Part. Opt. Part. Opt. Part. Opt. Part. Opt. Part. Opt.
CityGS [17] 188 61 77 51 118 79 97 78 71 55
VastGS [15] 51 89 4 44 10 59 8 68 20 52
Hier-GS [12] 9.8 119 12.7 101 8.0 137 4.4 130 7.6 120
Ours 1 67 <1 52 <1 78 <1 63 <1 59
Table 2: Partitioning and optimization efficiency of different methods across various scenes, with time measured in minutes.
@@
Baseline 0.747 25.06 0.281
add C1 0.779 25.40 0.244
add C2 0.807 25.74 0.229
add C3 0.834 26.33 0.201
Full model 0.839 26.42 0.197
Table 3: Ablation results on Rubble scene.
Ablation
To comprehensively assess the impact of each component in our proposed method, we conducted an ablation study on the Rubble scene. Specifically, we evaluate the effectiveness of our approach by progressively incorporating key components: visibility-based masking (C1), hierarchical weighted image supervision (C2), dynamic gradient thresholding (C3), and finally, our full model. Table 3 presents the corresponding results, demonstrating the effectiveness of our proposed strategies.
@@ 5 Conclusion
In this paper, we present HUG, a novel Hierarchical Urban Gaussian Splatting method designed to reconstruct and render large-scale aerial urban scenes with high quality. Comprehensive experimental results demonstrating State-of-the-Art performance of our model, proves the effectiveness of our proposed visibility-based partitioning, hierarchical weighted image supervision, dynamic gradient threshold, and other proposals.
Despite our approach’s advantages, a few limitations should be addressed in future work. First, our method does not account for dynamic objects. This can be improved by incorporating lighting decoupling and appearance encoding. Second, the data partitioning process relies on the sparse point cloud generated by COLMAP [27], which may lead to issues in low-texture areas. Combining pose-free Gaussian methods [10] may help it. Finally, our method does not support integrated aerial-ground training and rendering. This functionality could be incorporated in future research to enable more comprehensive scene reconstruction.
@@
We provide a more detailed discussion on the settings of the compared methods, view partition results, and additional qualitative comparisons. A rendered video is also included for a more comprehensive comparison.
6 Settings of Compared Methods
We compared our method against several state-of-the-art baselines, including Mega-NeRF [33], 3DGS [11], CityGS [17], OctreeGS [26], VastGS [15], and Hier-GS [12]. The quantitative results for Mega-NeRF [33], 3DGS [11], and CityGS [17] are taken from the CityGS [17] paper. The quantitative results for OctreeGS [26] are sourced from the OctreeGS [26] paper. The results for Hier-GS [12] were obtained using the released source code with all default settings. The results for VastGS [15] are derived from the unofficial code available at https://github.com/kangpeilun/VastGaussian, with the same block number as CityGS [17] and ours. The qualitative results for CityGS [17] were obtained from the checkpoint provided by the authors.
@@ 7 Data Partitioning Results
Figure 6: Comparison of view partition results among our method, VastGS [15], and CityGS [17] across four datasets. The radial coordinate indicates the number of training images assigned to each block, with each vertex representing a specific block. As shown, our method consistently assigns fewer images on average, demonstrating a more efficient partition strategy.
To associate all training images with their respective blocks, we compute the total number of visible sparse points in each view. A threshold of $\tau_{p}=800$ is then applied to determine the assignment of each training image to specific blocks. In CityGS [17], a coarse global 3DGS [11] model must first be constructed, followed by multiple renderings of all images corresponding to the number of blocks. VastGS [15] introduces a more complex camera selection strategy, incorporating position-based data selection, visibility-based camera selection, and coverage-based point selection, resulting in a significantly higher number of assigned cameras per block. VastGS[15] has an average of over 1,000 views per block on the MatrixCity dataset, which exceeds GPU memory limits for a single block. This is why we do not report its numerical results on this scene. For visual comparisons, we avoid that specific block whenever possible. In contrast, our data partitioning strategy assigns fewer images to each block, reducing memory consumption during training and allowing more computational resources to be dedicated to refining the in-block reconstruction. Figure 6 visualizes the number of training images per block, illustrating that our method allocates fewer images on average while still achieving superior rendering performance, as demonstrated in Table 1.
Figure 7: More qualitative comparisons on MatrixCity dataset.